# Resource-Efficient Beam Prediction in mmWave Communications with Multimodal Realistic Simulation Framework

### v1

@@ II-B Problem statement
The beam prediction model $f(\cdot)$ is parameterized by a set of parameters $\Theta$. These parameters are learned from the training dataset $\{\mathcal{X},\boldsymbol{b}^{*}\}$ at the channel sampling interval $t$, which contains the sensing information along with the corresponding optimal beam patterns. Consequently, the optimization problem predicting the beam for the future channel sampling interval $t+1$ can be written as [9]:
@@ III Multimodal Realistic Simulation Framework for Sensing-aided Communication
We now present a realistic multimodal simulation framework designed to generate multimodal sensor data and perform wireless communication experiments in a detailed virtual urban environment. Fig. 3 illustrates the workflow of the proposed framework. We use the autonomous driving tool CARLA [39] to generate realistic sensing information, while MATLAB [40] is used for communication experiments. Next, we describe the overall workflow, including procedures for sensor data generation and digital environment reconstruction. We then discuss how the reconstructed environment supports multimodal sensor simulation and ray-tracing-based channel generation, ultimately producing the data required for cross-modal knowledge distillation.
@@ IV-A Multimodal preprocessing
[…]
@@ IV-C1 Conventional Knowledge Distillation:
Conventional KD aims to transfer knowledge from the teacher to the student network by minimizing the loss of distillation. The distillation loss is calculated based on the difference between the features of the teacher and student networks. To calculate this difference, the conventional KD methods use the Kullback-Leibler (KL) divergence, which is a statistical measure that quantifies how a feature distribution $\mathcal{F}_{\mathrm{mono}}$ of the student network differs from a feature distribution $\mathcal{F}_{\mathrm{multi}}$ of the teacher network. For two feature distributions $P$ and $Q$, the KL divergence is calculated as [41]:
@@ IV-C2 Relational Knowledge Distillation:
While conventional KD aligns the characteristics of the teacher and student element by element, relational knowledge distillation further captures pairwise relationships between data samples or feature embeddings. This is particularly relevant for cross‐modal transfers, where the teacher’s input space (multimodal) differs from the student’s (radar only). By distilling the relational structure, the student learns how the teacher organizes features in a manifold, even without direct access to the teacher’s extra modalities.
(a) RGB samples of camera sensors in the 2-Lane scenario.
(b) RGB samples of camera sensors in the 3-Lane scenario.
Figure 5: RGB samples of camera sensors by episode type.
@@
To extract deep features from various relationships, we adopted a method called DistilVPR [42]. To capture higher-order relationships between feature embeddings, DistilVPR computes pairwise distances or similarities in three distinct manifolds: Euclidean (flat), spherical (positive curvature), and hyperbolic (negative curvature). For any two feature vectors $\boldsymbol{t}_{i}$ and $\boldsymbol{t}_{j}$, we denote:
@@
[…]
@@
$L_{\mathrm{rel}}=\sum_{i,j=1}^{B}\Bigl{[}$ $\,d\bigl{(}r_{\mathrm{euc}}(t_{i},t_{j}),\,r_{\mathrm{euc}}(s_{i},s_{j})\bigr{)}$
@@
where $d(\cdot,\cdot)$ is a distance or divergence. This ensures that the student maintains the geometry of the higher order of the teacher, even if the absolute feature values differ because the teacher sees more modalities. Consequently, we can replace the conventional KD term in (9) with relational loss as:
$\mathcal{L}_{\mathrm{rkd}}=(1-\alpha)\,L_{\mathrm{focal}}+\alpha\sum_{l\in\{\mathrm{mid},\,\mathrm{end}\}}L_{\mathrm{rel}}\!\Bigl{(}f_{t}^{(l)},\,f_{s}^{(l)}\Bigr{)},$ (15)
Similarly the conventional KD loss function (9), $\alpha$ is a weight parameter that balances the importance of the original loss and the distillation loss. $f_{t}^{(l)}$ and $f_{s}^{(l)}$ are the teacher and student feature maps of the layer $l$.
V Simulation Results and Analysis
@@
where $N$ is the total number of test samples, and $\mathbb{I}(\cdot)$ is an indicator function that returns $1$ if the condition is true and 00 otherwise. A higher Top-$k$ value indicates that the model’s probability distribution consistently assigns high scores to the correct beam within its top predictions.
2) Mean RSS: Once the model selects $\hat{y}$ as the predicted beam index, the system applies the corresponding beamformer, resulting in a received signal strength $S(\hat{y})$. We define the mean RSS as the average RSS achieved over the test set as follows:
@@
$\mathrm{rank}\bigl{(}\hat{y}\bigr{)}=1+\sum_{b=1}^{B}\mathbb{I}\Bigl{(}S(b)>S(\hat{y})\Bigr{)},$ (18)
@@
which ranges from 00 to $1$. The MPR over the test set can be written as:
@@
Methods Scenarios
2-Lane 3-Lane All
MPR (%) RSS (dBm) MPR (%) RSS (dBm) MPR (%) RSS (dBm)
@@
KD-mid 83.776 -82.561 78.293 -86.615 82.019 -84.538
KD-end 84.426 -82.322 79.503 -86.551 82.783 -84.412
KD-both 86.329 -81.399 79.860 -86.155 82.336 -84.209
RKD-mid 86.998 -81.266 79.525 -86.395 83.551 -84.006
RKD-end 85.844 -81.627 79.707 -86.323 83.463 -84.158
RKD-both 87.210 -81.084 80.101 -85.708 83.658 -83.603
@@
TABLE III: Comparative analysis of rank accuracy across different network size (6M, 13M, 34M) of student models.
Methods MPR (%) -
WithoutKD (6M) 83.272
RKD-both (6M) 86.772 3.500 ($\uparrow$)
WithoutKD (13M) 84.200
RKD-both (13M) 87.210 3.011 ($\uparrow$)
WithoutKD (34M) 86.364
RKD-both (34M) 88.007 1.643 ($\uparrow$)
TABLE IV: Comparison of beam prediction results under domain-same (2-Lane) and domain-shift (3-Lane to 2-Lane) scenarios.
@@
Teacher 92.172 -77.753 85.4962 -81.7826 84.3623 -84.1484 84.9292 -82.9655
WithoutKD 84.1999 -82.4066 53.3076 -94.5673 78.2879 -86.8027 65.7978 -90.6850
RKD 85.3297 -81.8894 55.0370 -93.9985 78.6399 -86.7685 66.8384 -90.3835
KD 84.8495 -82.1217 54.9017 -94.0079 77.7960 -86.9440 66.3488 -90.4760
@@
Table II summarizes the performance of cross-modal knowledge distillation for beam prediction. The Teacher network, a transformer-based model trained on multimodal sensing input (LiDAR, radar, GPS, and RGB), achieves the highest-rank accuracies of $92.17\%$, $84.36\%$, and $88.41\%$ in the scenarios tested. The 3-Lane scenario, having more complex multipath conditions, naturally results in lower accuracy across all models (Teacher’s MPR drops by approximately 8 points compared to the 2-Lane scenario), highlighting the added difficulty. The WithoutKD model is a radar-only MLP network without any distillation. The WithoutKD model reaches $84.12\%$, $78.29\%$, and $82.02\%$, demonstrating a lower performance than Teacher. This gap underscores the value of multimodal sensor data in guiding beam selection. We apply distillation at two levels of the network: a mid-level feature layer (the intermediate latent features, e.g. after the encoder or Transformer module) and the end layer (the final output logits before the softmax). KD-mid means applying conventional KD on an intermediate feature representation, KD-end to the output probabilities, and KD-both to both. A notable finding is that the proposed RKD-both method, a relational knowledge distillation strategy, achieves $87.21\%$, $80.10\%$, and $83.66\%$ while retaining a similar compact MLP architecture as WithoutKD. Compared to the KD-both approach with scores of $87.00\%$, $79.53\%$, and $80.64\%$, RKD-both offers an additional accuracy gain. These results indicate that even a radar-only student network can achieve high accuracy if it benefits from cross-modal distillation of teacher feature relationships. In practical terms, this finding suggests that a small, cost-effective network, limited to radar sensing for real-time inference, can still approximate the performance of a more complex multimodal teacher.
Figure 8 provides additional insight into these methods by illustrating the Top-$k$ performance, a commonly used metric in beam prediction scenarios. When $k=1$, RKD-both exhibits a considerably higher precision than WithoutKD, which is closely aligned with the rank precision reported in Table III. This consistency across different evaluation metrics (rank accuracy versus top-$k$) further validates that knowledge distillation with relational features improves overall beam prediction and ensures that the model’s most confident predictions are more frequently correct.
@@
Table III compares the performance of CRKD in different model sizes for the student network. The results indicate that as the MLP architecture of the student grows, the overall precision increases both for the baseline approach (WithoutKD) and the relational distillation approach (RKD-both). However, the relative improvement over KD is more pronounced in smaller student networks. In other words, while RKD-both achieves greater absolute accuracy with larger models, the performance gap between WithoutKD and RKD-both is notably greater when the student model is small. This suggests that knowledge distillation can be especially advantageous in resource-constrained scenarios, where the student network is required to maintain a low parameter count while striving to approximate the performance of a significantly more complex multimodal teacher. Consequently, compact radar-only models can benefit substantially from cross-modal distillation, realizing stronger beam prediction capabilities without incurring the computational overhead that larger networks demand.
@@
In addition to evaluating performance under matched training and testing domains, we also investigate a domain-shift scenario to assess how well knowledge distillation withstands different environmental conditions. Specifically, the teacher network is trained on a 2-Lane dataset and then used to distill knowledge into a student network trained on a 3-Lane dataset. After this cross-domain distillation, the student model is evaluated in both the 2-Lane and 3-Lane test sets. Table IV presents the beam prediction performance in both domain-matched (2-Lane) and domain-shifted scenarios, where the student model is trained on 3-Lane data and tested on 2-Lane and 3-Lane environments. The teacher model, trained on 2-Lane data, is also evaluated on both test sets for comparison. The Domain Shift setting specifically examines how well the student generalizes when trained and tested in different environments. The results show a noticeable drop in student performance (of up to $29.95\%$) under domain shift, especially in the 2-Lane test set, indicating limited generalization capacity. This performance degradation is more severe for the radar-only student due to its lower model capacity and restricted sensor input. In contrast, the multimodal teacher, with richer inputs and a transformer-based architecture, exhibits more stable performance across domains. These findings highlight the vulnerability of compact single-modal models to domain mismatch and underscore the importance of cross-domain robustness in real-world deployments.

### v2

@@ II-B Problem statement
The beam prediction model $f(\cdot)$ is parameterized by a set of parameters $\Theta$. These parameters are learned from the training dataset $\{\mathcal{X},\boldsymbol{b}^{*}\}$ at the channel sampling interval $t$, which contains the sensing information along with the corresponding optimal beam patterns. Consequently, the optimization problem predicting the beam for the future channel sampling interval $t+1$ can be written as [jiang2022lidar]:
@@ III Multimodal Realistic Simulation Framework for Sensing-aided Communication
We now present a realistic multimodal simulation framework designed to generate multimodal sensor data and perform wireless communication experiments in a detailed virtual urban environment. Fig. 3 illustrates the workflow of the proposed framework. We use the autonomous driving tool CARLA [Dosovitskiy17] to generate realistic sensing information, while MATLAB [MATLAB] is used for communication experiments. Next, we describe the overall workflow, including procedures for sensor data generation and digital environment reconstruction. We then discuss how the reconstructed environment supports multimodal sensor simulation and ray-tracing-based channel generation, ultimately producing the data required for cross-modal knowledge distillation.
@@ IV-A Multimodal preprocessing
  • LiDAR (Bird’s‐Eye View (BEV)): Raw LiDAR point clouds $\bigl[\text{x},\text{y},\text{z},\text{intensity}\bigr]$ are projected onto a 2D plane to form a BEV representation. This transformation highlights object occupancy and relative positioning in a top‐down image layout, filtering out ground reflections, and simplifying 3D geometry into 2D grids.
  • Radar (Highest Point Sampling (HPS)): To handle radar data $\bigl[\text{velocity},\,\text{azimuth},\,\text{altitude},\,\text{depth}\bigr]$, we apply HPS, which downsamples radar point clouds to a fixed size, while ensuring consistency across different sampling intervals or noise conditions.
@@ IV-C1 Conventional Knowledge Distillation:
Conventional KD aims to transfer knowledge from the teacher to the student network by minimizing the loss of distillation. The distillation loss is calculated based on the difference between the features of the teacher and student networks. To calculate this difference, the conventional KD methods use the Kullback-Leibler (KL) divergence, which is a statistical measure that quantifies how a feature distribution $\mathcal{F}_{\mathrm{mono}}$ of the student network differs from a feature distribution $\mathcal{F}_{\mathrm{multi}}$ of the teacher network. For two feature distributions $P$ and $Q$, the KL divergence is calculated as [kullback1951information]:
@@
(a) Manifold-Based Relationship.
(b) Beam Steering based Relationship.
Figure 5: Relational Knowledge Distillation with different relationships.
@@ IV-C2 Relational Knowledge Distillation:
While conventional KD focuses on aligning outputs or feature representations on a per-sample basis, relational knowledge distillation (RKD) goes a step further by preserving pairwise relationships between samples in the feature space. This approach is especially beneficial in cross-modal knowledge distillation, where the teacher model operates on richer multimodal data while the student has access to only a single modality (i.e., radar). In our framework, we implement two methods of relational knowledge distillation: (a) a manifold-based loss that aligns the latent feature geometry between teacher and student, and (b) a novel beam steering-based loss that captures the relational structure in the complex-valued beamspace at the output level. These complementary losses enable the student to better mimic the teacher’s internal representations and output behaviors under sensor constraints.
@@
To extract deep features from various relationships of the latent layer, we adopted a method called DistilVPR [wang2024distilvpr]. To capture higher-order relationships between feature embeddings, DistilVPR computes pairwise distances or similarities in three distinct manifolds: Euclidean (flat), spherical (positive curvature), and hyperbolic (negative curvature). For any two feature vectors $\boldsymbol{t}_{i}$ and $\boldsymbol{t}_{j}$, we denote:
@@
where $\langle\cdot,\cdot\rangle$ represents the inner (dot) product. In practice, one may use $1-r_{\mathrm{cos}}$ as a distance-like measure. Cosine-based relationships are useful when the magnitude of vectors is less important than their direction, a scenario common in classification or retrieval tasks. We incorporate a hyperbolic measure using the Poincaré ball model to capture hierarchical or tree-like structures and negative curvature. When $\boldsymbol{t}_{i},\boldsymbol{t}_{j}\in\mathbb{R}^{D}$, we can project each vector onto the Poincare space $\mathcal{D}_{c}^{D}$ (with curvature parameter $c>0$) via an exponential mapping as follows:
@@
$L_{\text{latent-rel}}=\sum_{i,j=1}^{B}\Bigl[$ $\,d\bigl(r_{\mathrm{euc}}(t_{i},t_{j}),\,r_{\mathrm{euc}}(s_{i},s_{j})\bigr)$
@@
where $d(\cdot,\cdot)$ is a distance or divergence. This ensures that the student maintains the geometry of the higher order of the teacher, even if the absolute feature values differ because the teacher sees more modalities.
Beam Steering based Relationship Measures:
In addition to the latent space structure, we propose a novel relational distillation loss in the output domain by leveraging the domain-specific properties of mmWave beamforming. Each predicted beam index corresponds to a complex-valued steering vector, which defines the phase alignment of signals transmitted across antenna elements. These vectors capture the directionality of beamforming and inherently encode spatial relationships between users or signal paths. Since steering vectors represent spatial directionality, aligning their pairwise similarities enables the student to learn beam-space relationships that are critical for mmWave propagation, where multipath components and direction selectivity play a dominant role in determining communication performance.
To preserve the relational structure of the beamspace, we compute a similarity matrix using the Hermitian inner product between normalized steering vectors. For two predicted beam indices $\hat{b}_{i}$ and $\hat{b}_{j}$ from the teacher model, their similarity is defined as:
[…]
where $\mathbf{w}(\hat{b}_{i})\in\mathbb{C}^{N}$ is the steering vector for the beam index $\hat{b}_{i}$, denotes the Hermitian (conjugate transpose), and $N$ is the number of antenna elements. The same operation is performed for the predicted beams of the student to yield $s_{i,j}^{(S)}$.
We then define a beamspace relational loss to align the structural similarities of beam directions between teacher and student:
$L_{\text{output-rel}}=\sum_{i,j=1}^{B}\left(s_{i,j}^{(T)}-s_{i,j}^{(S)}\right)^{2},$ (17)
where $B$ is the batch size. This loss ensures that, even without direct access to multimodal inputs, the radar-only student can mimic the teacher’s understanding of directional similarity across samples. In scenarios with domain shifts or multiple strong paths, preserving this angular relational structure is critical for robust beam prediction. To the best of our knowledge, this is the first application of relational knowledge distillation based on complex-valued beam steering vector similarities for mmWave communication. While prior work like DistilVPR focuses on latent space geometry, our proposed loss explicitly leverages directional structure at the output level, a domain-specific innovation tailored to beamforming systems under sensor constraints. Consequently, we can replace the conventional KD term in (10) with relational loss as:
$\mathcal{L}_{\mathrm{rkd}}=(1-\alpha)\,L_{\mathrm{focal}}+\alpha(L_{\text{latent-rel}}+L_{\text{output-rel}}),$ (18)
Similar to the conventional KD loss function (10), the parameter $\alpha$ balances the importance between the original supervised loss and the relational distillation losses.
(a) RGB samples of camera sensors in the 2-Lane scenario.
(b) RGB samples of camera sensors in the 3-Lane scenario.
Figure 6: RGB samples of camera sensors by episode type.
@@ V Simulation Results and Analysis
where $N$ is the total number of test samples, and $\mathbb{I}(\cdot)$ is an indicator function that returns $1$ if the condition is true and 0 otherwise. A higher Top-$k$ value indicates that the model’s probability distribution consistently assigns high scores to the correct beam within its top predictions.
2) Mean RSS: Once the model selects $\hat{y}$ as the predicted beam index, the system applies the corresponding beamformer, resulting in a received signal strength S(y^)S($\hat${y}). We define the mean RSS as the average RSS achieved over the test set as follows:
@@
[…]
@@
which ranges from 0 to $1$. The MPR over the test set can be written as:
@@ V-A Dataset Analysis
Scenarios
Methods 2-Lane 3-Lane All
Latent Output MPR (%) RSS (dBm) MPR (%) RSS (dBm) MPR (%) RSS (dBm)
@@
KLdiv - 83.776 -82.561 78.293 -86.615 82.019 -84.538
KD - KLdiv 84.426 -82.322 79.503 -86.551 82.783 -84.412
KLdiv KLdiv 86.329 -81.399 79.860 -86.155 82.336 -84.209
manifold - 86.998 -81.266 79.525 -86.395 83.551 -84.006
- manifold 85.844 -81.627 79.707 -86.323 83.463 -84.158
RKD manifold manifold 87.210 -81.084 80.101 -85.708 83.658 -83.603
- beamStr 87.185 -80.690 80.902 -85.624 83.276 -83.979
manifold beamStr 87.502 -80.610 81.116 -85.534 83.972 -83.526
@@
TABLE III: Comparison of MPR (%) with and without RKD across different student model sizes.
Model Size Without KD With RKD Gain
Student-S (6M) 83.27 86.77 +3.50
Student-M (13M) 84.20 87.50 +3.30
Student-L (34M) 86.36 88.45 +2.09
TABLE IV: Comparison of model complexity and inference latency across platforms.
Model #Params FLOPs GPU CPU1 CPU2
(ms) (ms) (ms)
Teacher 105.6M 20.3G 24.52 30.03 946.18
Student-L 34.2M 34.2M 0.66 0.99 7.99
Student-M 13.6M 13.6M 0.73 1.10 4.93
Student-S 6.8M 6.8M 0.59 1.04 2.21
TABLE V: Comparison of beam prediction results under domain-same (2-Lane) and domain-shift (3-Lane to 2-Lane) scenarios.
@@
Teacher 92.172 -77.753 85.496 -81.782 84.362 -84.148 84.929 -82.965
WithoutKD 84.199 -82.406 53.307 -94.567 78.287 -86.802 65.797 -90.685
sRKD 87.502 -80.610 54.781 -94.088 81.115 -85.625 67.802 -90.177
mRKD 85.329 -81.889 55.037 -93.998 80.101 -85.708 66.838 -90.383
KD 84.849 -82.121 54.901 -94.007 79.860 -86.155 66.348 -90.476
V-B Beam Prediction Accuracy
@@
Table II summarizes the performance of cross-modal knowledge distillation for beam prediction. The Teacher network, a transformer-based model trained on multimodal sensing input (LiDAR, radar, GPS, and RGB), achieves the highest-rank accuracies of $92.17\%$, $84.36\%$, and $88.41\%$ in the scenarios tested. The 3-Lane scenario, having more complex multipath conditions, naturally results in lower accuracy across all models (Teacher’s MPR drops by approximately 8 points compared to the 2-Lane scenario), highlighting the added difficulty. The WithoutKD model is a radar-only MLP network without any distillation. The WithoutKD model reaches $84.12\%$, $78.29\%$, and $82.02\%$, demonstrating a lower performance than Teacher. This gap underscores the value of multimodal sensor data in guiding beam selection. We apply distillation at two levels of the network: the latent layer (the latent features, e.g. after the encoder or Transformer module) and the output layer (the final output logits before the softmax). Among the standard KD baselines (using KL divergence), KD (KLdiv+KLdiv) reaches up to $86.32\%$, $79.86\%$, and $82.34\%$ but still underperforms compared to the RKD approaches. In the RKD section, various combinations of latent- and output-level relation transfer are explored: RKD (manifold+beamStr) achieves the highest student performance, with an MPR of $87.50\%$ on the 2-Lane and $83.97\%$ in the overall dataset. This configuration also yields the best RSS values (e.g. 80.61-80.61 dBm in 2-Lane), confirming its superior beam selection accuracy. The improvement of RKD over traditional KD methods illustrates the importance of transferring structured relationships, especially those based on beamforming-specific steering vector similarities (beamStr). Overall, these results indicate that even a radar-only student network can achieve high accuracy if it benefits from cross-modal distillation of teacher feature relationships. In practical terms, this finding suggests that a small, cost-effective network, limited to radar sensing for real-time inference, can still approximate the performance of a more complex multimodal teacher.
Fig. 9 illustrates the Top-K prediction accuracy ($K=1,5,10$) across five different models trained using all datasets. The baseline WithoutKD (radar-only MLP without distillation) shows the lowest performance across all K values, reflecting the limitations of lightweight monomodal learning. The KD model refers to conventional knowledge distillation using KL divergence at both the latent and output layers (KD (KLdiv+KLdiv)). The KD improves the accuracy of Top-1 and Top-10 compared to baseline, but remains less effective in capturing spatial beam relationships. The mRKD model represents relational knowledge distillation applied to both latent features (RKD (manifold+manifold), while sRKD applies RKD jointly on latent and output-level beam steering representations (RKD (manifold+beamStr)). Among the two, sRKD achieves superior performance, especially in Top-1 and Top-5 metrics, highlighting the benefit of beam-aware similarity modeling in the output space. The Teacher model, trained on full multimodal input with transformer architecture, achieves the highest overall accuracy, serving as an upper bound. However, sRKD narrows the performance gap in Top-10 accuracy while maintaining a lightweight radar-only architecture, demonstrating the effectiveness of beam-aware relational distillation under resource constraints.
@@
Table III compares the performance of RKD in different model sizes for the student network. The results indicate that as the MLP architecture of the student grows, the overall precision increases both for the baseline approach (WithoutKD) and the relational distillation approach (RKD). However, the relative improvement over KD is more pronounced in smaller student networks. In other words, while RKD achieves greater absolute accuracy with larger models, the performance gap between WithoutKD and RKD is notably greater when the student model is small. This suggests that knowledge distillation can be especially advantageous in resource-constrained scenarios, where the student network is required to maintain a low parameter count while striving to approximate the performance of a significantly more complex multimodal teacher. Consequently, compact radar-only models can benefit substantially from cross-modal distillation, realizing stronger beam prediction capabilities without incurring the computational overhead that larger networks demand.
V-C Resource Efficiency Analysis
Table IV presents a comparison of model size, FLOPs, and inference latency for teacher and three student variants on three hardware platforms: a GPU (NVIDIA RTX 4070), a desktop CPU (AMD Ryzen 7 7800X3D @ 4.2GHz) and a mobile CPU (Intel i7-10510U @ 1.8GHz). While the teacher model, based on a Transformer architecture, incurs high computational cost (20.3G FLOPs) and latency up to 946.18ms on mobile CPUs, the student models are designed with lightweight MLPs. As a result, their FLOPs closely match their parameter counts (e.g. Student-S: 6.8M parameters $\approx$ 6.8M FLOPs), enabling significant speed up during inference. In particular, the smallest student model achieves 0.59ms on GPU and 1.04ms on CPU1, satisfying the sub-1ms latency target for 5G/6G beamforming, especially under ultra-reliable low-latency communication (URLLC) requirements. Moreover, given that standard 5G NR beam sweeping takes approximately 20ms, our method provides over 30× faster inference with competitive accuracy. This indicates the practical deployability of our method for real-time mmWave beam prediction on edge and mobile devices, without compromising performance.
V-D Limitations under Domain Shift and Future Directions
@@
In addition to evaluating performance under matched training and testing domains, we also investigate a domain-shift scenario to assess how well knowledge distillation withstands different environmental conditions. Specifically, the teacher network is trained on a 2-Lane dataset and then used to distill knowledge into a student network trained on a 3-Lane dataset. After this cross-domain distillation, the student model is evaluated in both the 2-Lane and 3-Lane test sets. Table V presents the beam prediction performance in both domain-matched (2-Lane) and domain-shifted scenarios, where the student model is trained on 3-Lane data and tested on 2-Lane and 3-Lane environments. The teacher model, trained on 2-Lane data, is also evaluated on both test sets for comparison. The Domain Shift setting specifically examines how well the student generalizes when trained and tested in different environments. The results show a noticeable drop in student performance (of up to $32.72\%$) under domain shift, especially in the 2-Lane test set, indicating limited generalization capacity. This performance degradation is more severe for the radar-only student due to its lower model capacity and restricted sensor input. In contrast, the multimodal teacher, with richer inputs and a transformer-based architecture, exhibits more stable performance across domains. These findings highlight the vulnerability of compact single-modal models to domain mismatch and underscore the importance of cross-domain robustness in real-world deployments.